8. Each of the main cables supporting a suspension bridge can be roughly modeled as a one- dimensional string of length 300m with a mass of 5x106kg, placed under tension of 10 MN with 1.0 x 107 N. Suddenly, an earthquake creates a transverse wave in the cable an amplitude of 1.60m. a. What is the wave speed? b. If the earthquake drives the wave with a frequency 1.3Hz, what is the wavelength? of c. What is the wave number? d. What is the angular frequency? e. How much power is being delivered to the skyscraper by the earthquake?
Answer #1

a)

wave speed v = sqrt(T/u)

T = tension force = 10^7 N


u = mass/length = 7.5*10^6/300 = 25000 kg /m


speed v = sqrt(10^7/25000) = 20 m/s

---------

(b)wavelength = speed/ frequency


wavelength = 20/(1.3) = 15.4 m


-------------

(c)

wave number K = 2*pi/wavelength

K = 2*pi/15.4 = 0.41 m^-1

------


(d)

angular frequency w = v*k = 2*pi*f = 8.2 rad/s


-------

(e)


power P = (1/2)*u*w^2*A^2*v

P = (1/2)*25000*8.2^2*1.6^2*20

P = 43033600 W